the volume of acetylene in a tank with a tight-fitting piston is 5.70 cubic feet. the gage pressure is 28.2 psi. the volume of the tank is increased until the new gage pressure is 20.0 psi. what is the new volume of the acetylene in the tank? assume that the temperature remains the same as the volume increased and that there is standard atmospheric pressure.


1
Expert's answer
2021-04-08T09:00:02-0400

We have given that

"P_1 = 28.2" psi

"P_2 = 20" psi

"V_1 = 5.70" cubic feet

"V_2 = x"

Using the equation of continuity in Fluid,


"P_1V_1 = P_2V_2"


"28.2\times 5.70 = 20\times V_2"


"V_2 = \dfrac{160.74}{20}"


"V_2 = 8.037" cubic feet.